# Rotating Einstein–Maxwell black holes in odd dimensions

### v1

@@ 1 Introduction
Over the last several decades there has been a remarkable progress in understanding of black holes in higher dimensions and in applying them to studies of quantum gravity [1] and strongly coupled systems [2, 3]. In the absence of rotation, one can easily generalize the Reissner–Nordstrom black holes to arbitrary dimensions by adding charges to the Schwarzschild–Tangherlini solutions. In the rotating case, even the neutral solutions turned out to be rather nontrivial, and they are given by the Myers–Perry geometry [4] and its generalization to the system with cosmological constant [5]. Charged versions of the Myers–Perry geometry can be constructed using the U–duality techniques developed in [6]: they involve adding a torus to the system and performing various string dualities. In the presence of the cosmological constant, the procedure of adding charges is much less straightforward, and it had been applied only on the case–by–case basis by embedding the systems into gauged supergravities [^3] Examples of charged solutions in gauged supergravities in various dimensions include [3, 7, 8, 9, 10, 11]. Beyond gauged supergravities, very few charged asymptically–AdS geometries have been constructed [12].. Both the duality techniques and gauged supergravities produce solutions which have various scalars in addition to gauge fields, so even in four dimensions the geometries don’t reduce to the Kerr-Newman (KN) black holes.
@@
Given the importance of the Kerr-Newman geometry, it is natural to ask for higher dimensional generalization of this solution supported only by a gauge field. Unfortunately, numerous attempts to find rotating solutions of the Einstein–Maxwell theory beyond four dimensions have failed [13], but in article [14] a very close cousin of the KN solution was constructed in five dimensions. Specifically, the authors of [14] solved the Einstein’s equations with Maxwell sources, but they modified the equations for the gauge fields by adding a topological term to the action. We will extend this logic to all odd dimensions by constructing rotating solutions of Einstein’s equation with Maxwell sources in the presence of a negative cosmological constant. To satisfy equation of motion for the gauge field, we will add some topological terms to the action, and beyond five dimensions such terms will involve non–dynamical auxiliary fields. The resulting solutions are probably the closest analytical counterpart of the KN black holes in higher dimensions. In this article we focus on odd–dimensional solutions with equal rotations since in this case the symmetries of the problem lead to a relatively simple ansatz where most metric components are functions of only one coordinate. In section 2 we will comment on challenges associated with extending this ansatz to non–equal rotations.
@@
One of the most remarkable properties of the Myers–Perry and Gibbons–Lu–Page–Pope geometries [4, 5] is full integrability of dynamical equations for all fields on these backgrounds. As shown in [15], these geometries admit a family of Killing–Yano tensors (KYT) which guarantee separability of equations for scalars and spinors [^4] In [16] the construction of KYTs was extended to charged geometries obtained from the Myers–Perry solutions by applying string dualities.. The equations for vector fields and higher forms were also shown to be separable [17, 18, 19]. This raises a question about persistence of such integrable structures for the charged geometries constructed in section 2 of our article. In section 4 we demonstrate that the equation for a charged scalar field remains fully separable in all dimensions, while the Killing–Yano tensors seem to disappear [^5] Such disappearance has been already noticed for the five–dimensional CCLP black holes constructed in [14], although in that case one could define a KYT with torsion [20]. Unfortunately, the procedure for introduction of torsion from the Chern–Simons coupling seems to be specific to five dimensions..
@@
This paper has the following organization. In section 2 we construct charged rotating geometries in all odd dimensions by proposing an ansatz for the metric and solving the Einstein’s equations sourced by a gauge field. We also propose the Chern–Simons coupling to additional auxiliary fields which ensure that equations for the gauge field are satisfied as well. This Chern–Simons coupling is similar to the one encountered in the minimal five–dimensional supergravity in [14], although in that case no fields in addition to the vector were required. In section 3, we analyze various physical properties of the new solutions, such as their entropies, temperatures, and charges. In section 4 we study dynamical excitations of the new solutions and show that the charged Klein–Gordon equation is fully separable in all odd dimensions. Some technical details are presented in appendices. Specifically, Appendix A reviews the Chong–Cvetic–Lu–Pope (CCLP) solution of the minimal five–dimensional dimensional supergravity which serves as an inspiration for our construction. In the same appendix we also demonstrate full separability of the charged Klein–Gordon equation in the CCLP geometry to connect to our discussion in section 4. Appendix B reviews the neutral rotating black holes in all odd dimensions and special frames that play important role in ensuring separability of various dynamical equations. In the same appendix we also take a singular limit of the known frames to cover the case of equal rotations relevant for sections 2 and 4. Appendix C outlines the general procedure for computing surface gravity of higher dimensional black holes and applies it to our new solutions. Appendix D provides some technical details supporting the analysis presented in section 4.
@@ 2 Generalization of the KN–AdS black hole
The main goal of this article is to generalize the Kerr-Newman geometry to higher dimensions, i.e., to construct solutions of the Einstein’s equations sourced only by a gauge field. Previous attempts to solve this problem have failed [13] due to a clash between Einstein and Maxwell equations. This clash is avoided in the minimal five–dimensional supergravity, which keeps the Einstein’s equations from the Einstein–Maxwell theory but modifies the equations for the gauge field by adding a topological term to the action. A close five–dimensional counterpart of the Kerr-Newman solution in this theory was found in [14]. To generalize this construction to higher dimensions, we consider the system governed by the action
@@
Inspired by the charged five–dimensional solution [14] reviewed in the Appendix A and by the neutral rotating black holes in arbitrary dimensions [4, 5] reviewed in the Appendix B, we consider an ansatz for the metric and the gauge field in $d=2s+3$ dimensions [^6] To simplify various formulas appearing in this article, we will interchangeably parameters $s$ and $n$ related to the dimensions $d$ by $d=2n+1=2s+3$.
@@
which matches (2.7) only for $n=2$. Therefore, the standard Chern–Simons term (2.8) can’t be used to generate equations (2.7) beyond five dimensions, so we have to introduce some additional auxiliary fields. Specifically, we consider the action
@@
where $f_{3}$ is an undetermined function, and $\nu$ is an one–form with undetermined coefficients $g_{i}$
@@
Unfortunately, for unequal rotations in $d>5$, we were not able to extend (2)–(2.18) to the full charged solution at all orders in $q$.
To summarize, we have demonstrated that the geometry (2)–(2.6) solves equations of motion coming from the action (2.9). The Einstein’s equations (2.3) have sources only from the Maxwell terms, just as they do for the Kerr–Newman metric in four dimensions, while the generalized Maxwell’s equations (2.7) contain sources from the auxiliary fields $(G^{(n)},{\bar{G}}^{(n)})$. In the five–dimensional case, the geometry (2)–(2.6) reduces to the solution found in [14]. Interestingly, in contrast to the four–dimensional Kerr–Newman case, where metric contains only even powers of charge $q$, the solution (2) has both even and odd powers. This phenomenon already appeared in the five dimensional case studied in [14].
@@ 3 Physical parameters of the solution
In the previous section we constructed the charged rotating geometry (2)–(2.6) in all odd dimensions. Let us now analyze various conserved quantities associated with this solution.
@@ 4 Separation of the Klein–Gordon equation
In this section we will analyze excitations of the geometry (2). First, we recall that in the neutral case, this solution has a family of Killing–Yano tensors which guarantee separation of variables in equations for scalar and spinor fields [15, 21]. The equations for Maxwell and higher form bosonic fields are also separable [17, 18, 19]. As we will show below, introduction of charge destroys the Killing–Yano tensors, but the equation for the scalar field still remains fully integrable.
We begin with recalling that equations for excitations of the neutral solution with multiple rotations become separable in ellipsioidal coordinates [15]. In the Appendix B we take the degenerate limit of these coordinates which is necessary for describing equal rotations. As demonstrated in that appendix, such degenerate coordinates $(r,y_{1},\dots,y_{n-1})$ are given by solving relations
@@ 5 Discussion
In this article we focused on black holes which have equal rotation parameters in $n$ planes, and it would be very interesting to extend our results to solutions with several independent rotations. As discussed in section 2, unfortunately such an extension would require a significant modification of the ansatz, and it not clear that the relevant solution can be obtained in a closed analytical form. Although in this article we have focused on black holes in odd dimensions, it appears that charged geometries in even dimensions greater than four lead to the same difficulties as the case of unequal rotations. The situation might be similar to the one encountered in higher dimensional Einstein–Maxwell equations without topological couplings: while it is clear that charged rotating solutions exist, all attempts to find them analytically have failed.
@@ Appendix A Charged solution in five dimensions
In this appendix we briefly summarize the geometries of the CCLP charged black holes in the minimal five–dimensional supergravity [14], which served as an inspiration for the construction presented in section 2, and discuss separation of variables in charged Klein–Gordon equation on the CCLP background.
@@
To connect to the discussion presented in section 4, we will now analyze the Klein–Gordon equation in the geometry (A.2). For the neutral scalar field, such analysis was done in [22], and here we will extend these results to the charged field, and more importantly, we will introduce notation that can be generalized to higher dimensions discussed in section 4.
@@
For the neutral scalar (i.e., for $e=0$) these equations are equivalent to those found in [22]. If two rotation parameters are equal, $a_{1}=a_{2}$, then equations (A)–(A) reduce to the special case of (4) for $d=5$.
@@ Appendix B Neutral rotating black holes and their symmetries
In this appendix we review the metrics for neutral rotating black holes in arbitrary dimensions. In the absence of the cosmological constant, such solutions were constructed by Myers and Perry in [4], and the AdS generalization was found by Gibbons–Lu–Page–Pope [5]. As demonstrated in [15], scalar and spinor equations on the GLPP backgrounds were separable, and an important role in ensuring this property was played by special frames. In this appendix we review construction of these frames for generic values of rotation parameters and take a singular limit to accommodate equal rotations discussed in sections 2 and 4.
We begin with recalling the GLPP metric with multiple rotations [5]
@@
As demonstrated in [15], the geometry (B.1) admits a Principal Conformal Killing–Yano tensor (PCKYT), which implies full separability of equations for various dynamical fields in the background (B.1). This PCKYT can be written in a compact form in terms of some special frames. To arrive at these frames, we begin with introducing ellipsoidal coordinates $(r,x_{1},\dots x_{n-1})$, where $x_{k}$ are defined by
@@
Then special frames for the metric (B.1) are defined by [15]
@@
As demonstrated in [15], the geometry (B.1) admits a family of Killing–Yano tensors of various ranks which can be written in a compact form
@@
The family (B.8) implies full integrability of dynamical equations for various fields in the background (B.1), and we refer to [15, 21] for details. In this article we are interested in the degenerate limit of the frames (B) with equal rotation parameters $a_{k}$.
@@ Appendix C Evaluation of surface gravity
We begin with recalling the definition of the surface gravity (3.6) in terms of the Killing vector (3.4). Using the Killing equation, we can rewrite (3.6) as
@@
where the metric $ds_{\perp}^{2}$ built from non–cyclic coordinates remains regular at the horizon. Direct evaluation of (C.1) in the coordinate system (C.2) leads to singularities, and to cure this problem one defines a new set of coordinates $(v,\psi^{a})$ by

### v2

@@ 1 Introduction
Over the last several decades there has been a remarkable progress in understanding of black holes in higher dimensions and in applying them to studies of quantum gravity [1, 2, 3, 4] and strongly coupled systems [5, 6, 7, 8, 9, 10, 11]. In the absence of rotation, one can easily generalize the Reissner–Nordstrom black hole to arbitrary dimensions by adding charges to the Schwarzschild–Tangherlini solution. In the rotating case, even the neutral solutions turned out to be rather nontrivial, and they are given by the Myers–Perry geometry [12] and its generalization to the system with cosmological constant [13, 14]. Charged versions of the Myers–Perry geometry can be constructed using the U–duality techniques developed in [15, 16, 17, 18, 19, 20]: they involve adding a torus to the system and performing various string dualities. In the presence of the cosmological constant, the procedure of adding charges is much less straightforward, and it had been applied only on the case–by–case basis by embedding systems into gauged supergravities [^3] Examples of charged solutions in gauged supergravities in various dimensions include [9, 10, 11, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34]. Beyond gauged supergravities, very few charged asymptically–AdS geometries have been constructed [35, 36, 37].. Both the duality techniques and gauged supergravities produce solutions which have various scalars in addition to gauge fields, so even in four dimensions the geometries do not reduce to the Kerr-Newman (KN) black holes.
@@
Given the importance of the Kerr-Newman geometry, it is natural to ask for a higher dimensional generalization of this solution supported only by a gauge field. Unfortunately, numerous attempts to find rotating solutions of the Einstein–Maxwell theory beyond four dimensions have failed [38, 39, 40, 41, 42], but in article [43] a very close cousin of the KN solution was constructed in five dimensions. Specifically, the authors of [43] solved the Einstein’s equations with Maxwell sources, but they modified the equations for the gauge fields by adding a topological term to the action. We will extend this logic to all odd dimensions by constructing rotating solutions of Einstein’s equation with Maxwell sources in the presence of a negative cosmological constant. To satisfy equation of motion for the gauge field, we will add some topological terms to the action, and beyond five dimensions such terms will involve non–dynamical auxiliary fields. The resulting solutions are probably the closest analytical counterpart of the KN black holes in higher dimensions. In this article we focus on odd–dimensional solutions with equal rotations since in this case the symmetries of the problem lead to a relatively simple ansatz where most metric components are functions of only one coordinate. In section 2 we will comment on challenges associated with extending this ansatz to non–equal rotations.
@@
One of the most remarkable properties of the Myers–Perry and Gibbons–Lu–Page–Pope geometries [12, 13, 14] is full integrability of dynamical equations for all fields on these backgrounds. As shown in [44, 45, 46, 47, 48], these geometries admit a family of Killing–Yano tensors (KYT) which guarantee separability of equations for scalars and spinors [^4] In [49] the construction of KYTs was extended to charged geometries obtained from the Myers–Perry solutions by applying string dualities.. The equations for vector fields and higher forms were also shown to be separable [50, 51, 52, 53, 54]. This raises a question about persistence of such integrable structures for the charged geometries constructed in section 2 of our article. In section 4 we demonstrate that the equation for a charged scalar field remains fully separable in all dimensions, while the Killing–Yano tensors seem to disappear [^5] Such disappearance has been already noticed for the five–dimensional CCLP black holes constructed in [43], although in that case one could define a KYT with torsion [55]. Unfortunately, the procedure for introduction of torsion from the Chern–Simons coupling seems to be specific to five dimensions..
@@
This paper has the following organization. In section 2 we construct charged rotating geometries in all odd dimensions by proposing an ansatz for the metric and solving the Einstein’s equations sourced by a gauge field. We also propose the Chern–Simons coupling to additional auxiliary fields which ensure that equations for the gauge field are satisfied as well. This Chern–Simons coupling is similar to the one encountered in the minimal five–dimensional supergravity in [43], although in that case no fields in addition to the vector were required. In section 3, we analyze various physical properties of the new solutions, such as their entropies, temperatures, and charges. In section 4 we study dynamical excitations of the new solutions and show that the charged Klein–Gordon equation is fully separable in all odd dimensions. Some technical details are presented in appendices. Specifically, Appendix A reviews the Chong–Cvetic–Lu–Pope (CCLP) solution of the minimal five–dimensional dimensional supergravity which serves as an inspiration for our construction. In the same appendix we also demonstrate full separability of the charged Klein–Gordon equation in the CCLP geometry to connect to our discussion in section 4. Appendix B reviews the neutral rotating black holes in all odd dimensions and special frames that play an important role in ensuring separability of various dynamical equations. In the same appendix we also take a singular limit of the known frames to cover the case of equal rotations relevant for sections 2 and 4. Appendix C outlines the general procedure for computing surface gravity of higher dimensional black holes and applies it to our new solutions. Appendix D provides some technical details supporting the analysis presented in section 4.
@@ 2 Generalization of the KN–AdS black hole
The main goal of this article is to generalize the Kerr-Newman geometry to higher dimensions, i.e., to construct solutions of the Einstein’s equations sourced only by a gauge field. Previous attempts to solve this problem have failed [38, 39, 40, 41, 42] due to a clash between the Einstein’s and Maxwell’s equations. This clash is avoided in the minimal five–dimensional supergravity, which keeps the Einstein’s equations from the Einstein–Maxwell theory but modifies the equations for the gauge field by adding a topological term to the action. A close five–dimensional counterpart of the Kerr-Newman solution in this theory was found in [43]. To generalize this construction to higher dimensions, we consider the system governed by the action
@@
Inspired by the charged five–dimensional solution [43] reviewed in the Appendix A and by the neutral rotating black holes in arbitrary dimensions [12, 13, 14] reviewed in the Appendix B, we consider an ansatz for the metric and the gauge field in $d=2s+3$ dimensions [^6] To simplify various formulas appearing in this article, we will interchangeably parameters $s$ and $n$ related to the number of dimensions $d$ by $d=2n+1=2s+3$.
@@
which matches (2.7) only for $n=2$. Therefore, the standard Chern–Simons term (2.8) cannot be used to generate equations (2.7) beyond five dimensions, so we have to introduce some additional auxiliary fields. Specifically, we consider the action
@@
where $f_{3}$ is an undetermined function, and $\nu$ is a one–form with undetermined coefficients $g_{i}$
@@
Unfortunately, for unequal rotations in $d>5$, we were not able to extend the geometry (2)–(2.18) to the full charged solution at all orders in $q$.
To summarize, we have demonstrated that the geometry (2)–(2.6) solves equations of motion coming from the action (2.9). The Einstein’s equations (2.3) have sources only from the Maxwell terms, just as they do for the Kerr–Newman metric in four dimensions, while the generalized Maxwell’s equations (2.10) contain sources from the auxiliary fields $(G^{(n)},{\bar{G}}^{(n)})$. In the five–dimensional case, the geometry (2)–(2.6) reduces to the solution found in [43]. Interestingly, in contrast to the four–dimensional Kerr–Newman case, where the metric contains only even powers of charge $q$, the solution (2) has both even and odd powers. This phenomenon already appeared in the five dimensional case studied in [43].
@@ 3 Physical parameters of the solution
In the previous section we constructed the charged rotating geometry (2)–(2.6) in all odd dimensions. Let us now analyze various physical properties of this solution.
@@ 4 Separation of the Klein–Gordon equation
In this section we will analyze excitations of the geometry (2). First, we recall that in the neutral case, this solution has a family of Killing–Yano tensors which guarantee separation of variables in equations for scalar and spinor fields [44, 45, 46, 47, 48, 56, 57] [^11] In the absence of the cosmological parameter $g$, separation of variables in the backgrounds of higher dimensional black holes was explored in [58, 59, 60, 61, 62]. The equations for Maxwell and higher form bosonic fields are also separable [50, 51, 52, 53, 54]. As we will show below, introduction of charge destroys the Killing–Yano tensors, but the equation for the scalar field still remains fully integrable.
We begin with recalling that equations for excitations of the neutral solution with multiple rotations become separable in ellipsioidal coordinates [44, 45, 46, 47, 48]. In the Appendix B we take the degenerate limit of these coordinates which is necessary for describing equal rotations. As demonstrated in that appendix, such degenerate coordinates $(r,y_{1},\dots,y_{n-1})$ are given by solving relations
@@ 5 Discussion
In this article we focused on black holes which have equal rotation parameters in $n$ planes, and it would be very interesting to extend our results to solutions with several independent rotations. As discussed in section 2, unfortunately such an extension would require a significant modification of the ansatz, and it is not clear whether the relevant solution can be obtained in a closed analytical form. Although in this article we have focused on black holes in odd dimensions, it appears that charged geometries in even dimensions greater than four lead to the same difficulties as the case of unequal rotations. The situation might be similar to the one encountered in higher dimensional Einstein–Maxwell equations without topological couplings: while it is clear that charged rotating solutions exist, all attempts to find them analytically have failed.
@@ Appendix A Charged solution in five dimensions
In this appendix we briefly summarize the geometries of the CCLP charged black holes in the minimal five–dimensional supergravity [43], which served as an inspiration for the construction presented in section 2, and discuss separation of variables in charged Klein–Gordon equation on the CCLP backgrounds.
@@
To connect to the discussion presented in section 4, we will now analyze the Klein–Gordon equation in the geometry (A.2). For the neutral scalar field, such analysis was carried out in [63] [^13] Recently fluctuations of some vector and tensor fields on the CCLP background were studied as well [64]., and here we will extend these results to the charged field, and more importantly, we will introduce notation that can be generalized to higher dimensions discussed in section 4.
@@
For the neutral scalar (i.e., for $e=0$) these equations are equivalent to those found in [63]. If the two rotation parameters are equal, $a_{1}=a_{2}$, then equations (A)–(A) reduce to the special case of (4) for $d=5$.
@@ Appendix B Neutral rotating black holes and their symmetries
In this appendix we review the metrics for neutral rotating black holes in arbitrary dimensions. In the absence of the cosmological constant, such solutions were constructed by Myers and Perry in [12], and the AdS generalization was found by Gibbons–Lu–Page–Pope [13, 14]. As demonstrated in [44, 45, 46, 47, 48], scalar and spinor equations on the GLPP backgrounds were separable, and an important role in ensuring this property was played by special frames. In this appendix we review the construction of these frames for generic values of rotation parameters and take a singular limit to accommodate equal rotations discussed in sections 2 and 4.
We begin with recalling the GLPP metric with multiple rotations [13, 14]
@@
As demonstrated in [44, 45, 46, 47, 48], the geometry (B.1) admits a Principal Conformal Killing–Yano tensor (PCKYT), which implies full separability of equations for various dynamical fields in the background (B.1). This PCKYT can be written in a compact form in terms of some special frames. To arrive at these frames, we begin with introducing ellipsoidal coordinates $(r,x_{1},\dots x_{n-1})$, where $x_{k}$ are defined by
@@
Then the special frames for the metric (B.1) are defined by [44, 45, 46, 47, 48]
@@
As demonstrated in [44, 45, 46, 47, 48], the geometry (B.1) admits a family of Killing–Yano tensors of various ranks which can be written in a compact form
@@
The family (B.8) implies full integrability of dynamical equations for various fields in the background (B.1), and we refer to [44, 45, 46, 47, 48, 56, 57] for details. In this article we are interested in the degenerate limit of the frames (B) with equal rotation parameters $a_{k}$.
@@ Appendix C Evaluation of surface gravity
We begin with recalling the definition of the surface gravity (3.6) in terms of the Killing vector (3.3)–(3.4). Using the Killing equation, we can rewrite (3.6) as
@@
where the contribution $ds_{\perp}^{2}$ built from non–cyclic coordinates remains regular at the horizon. Direct evaluation of (C.1) in the coordinate system (C.2) leads to singularities, and to cure this problem one defines a new set of coordinates $(v,\psi^{a})$ by